OpenStudy (anonymous):

Help me Do THis Problem By Steps Factor 3a2-12a+ab-4b

OpenStudy (anonymous):

Factor by grouping: \[(3a^2-12a)+(ab-4b)\]

OpenStudy (ranga):

What is the greatest common factor that you can factor out of the first two terms above? ^^^ What is the greatest common factor that you can factor out of the last two terms above?

OpenStudy (ranga):

3 will go into 3a^2 and 3 will also go into 12a a will go into 3a^2 and a will also go into 12a So 3a can be factored out of the first two terms: 3a(a - 4) Do the same with ab - 4b

OpenStudy (ranga):

3a^2-12a+ab-4b = (3a^2-12a) + (ab-4b) = 3a(a - 4) + b(a-4) = (3a+b)(a-4) Those are the two factors.